### v1

@@ 4.1 Music Streaming
As noted above, measuring audience development can have a long time scale, so a short term proxy for this quality could be to measure to what extent an artist’s music is being recommended to receptive users. There are multiple ways to determine if a user is receptive [^3] For example, did the user listen to a second song by the artist, add their songs to a playlist, etc.?, but for the sake of example, let us assume that we can measure the number $n$ of non-audience listeners (that is, $u\notin A_{r}$) who were recommended a song by $r$ and then listened to the entire song. Given that musicians have very different numbers of fans, it might make sense to normalize by the size of the artist’s existing audience $A_{r}$: $m_{r}=n/|A_{r}|$.
As a metric shared with individual providers, a low score on $m_{r}$ might raise concerns for the artist relative to the recommender system. It would mean that few new listeners are being introduced to their music. For a superstar, this might not be an issue: many people know their music already, but for an emerging artist, it could indicate that the recommender is not working as it should. A higher $m_{r}$ score does not necessarily mean that their audience is growing, but it does mean that the recommender system is introducing their music to potential new fans. From the system stakeholder point of view, this score could also be aggregated across all providers to understand audience building across the platform’s stable of artists. Its distribution might also be relevant in terms of fairness: are some types of artists better able to build audiences on the platform than others?
@@ 4.2 Education
Figure 3: Stakeholder relations for the education example
@@ 4.3 Human Resources
Figure 4: Stakeholder relations for the human resources example
@@
5 Concluding Remarks
A holistic understanding of recommender system operation requires considering the perspectives of multiple parties beyond the users receiving recommendations. This area of recommender systems evaluation is relatively underrepresented in the research literature, although, in commercial settings, such considerations have always been an element of recommender system development. Throughout our discussions in Sections 1 to 4, we have emphasized some of the reasons why this work is challenging to conduct and therefore has seen limited research attention.
In our narrative thus far, we have described general properties of multistakeholder recommendation, and methodological approaches to developing relevant metrics, and investigated three hypothetical examples of metric development targeted to different classes of stakeholders. In addressing core ingredients of multistakeholder recommender system evaluation, we hope to inspire reflection of the challenges and possible solutions. In the rest of this section, we discuss some salient challenges that we have identified.
5.1 Transparency / Explainability
@@
However, the types of evaluations that we discuss are different in that they may be of interest to parties who normally have no access to the workings of the recommender system. For example, the musical artists in our streaming example typically have very little insight into how the recommender system treats their content. Earlier, we noted that a given musical artist might seek to understand to what extent is the recommender system helping them build an audience. Such a metric could be shared with artists as a form of explanation to help them understand what the recommender system is doing.
Explanations in a multistakeholder context bring challenges different from explanations targeted only toward consumers. Firstly, different stakeholder groups have different explanatory needs that need to be identified. In the aforementioned example, the artist, and their listeners, have different explanatory needs. The next question is whether one can present different explanations to other stakeholders, or whether the explanation needs to be given to all parties. For example, should we explain item recommendations to individual users and audience building to artists separately? Or is there a single explanation that explains how we resolve the exposure of artists relative to how we weigh user preferences? In other words, if the requirement is to find and generate such a general explanation, then we need systems that can generate a meta-explanation on how tensions were resolved. A particular tension has already been observed, where consumers (depending on the context of recommendation) may prefer less transparency in explanations if it gives better privacy [54, 77]. Members of groups preferred not to disclose sensitive information to other group members, and consumers of advertisements did not want certain information to be used in explanations (or as a basis of recommendations!). We do not attempt to answer the question of how to generate these kinds of explanations (as this is out of scope), but note that provider-side transparency, let alone generalized explanations, are very little studied in the context of multistakeholder recommendation (evaluation).
@@ 5.2 Strategic / Adversarial Considerations
One likely reason that multistakeholder transparency has been little pursued in recommender systems research is the concern that such a facility might be used to enable undesirable adversarial behavior. A web search for the term “YouTube algorithm” yields thousands of hits from search engine optimization (SEO) firms and others advising creators about how to bend the algorithm to their will. Additional information given to providers may enhance their ability to manipulate the algorithm in ways that are not necessarily beneficial to recommendation consumers or the platform. An open research question is how to offer provider-side disclosure in a way that limits adversarial opportunities.
@@ 5.3 Governance
5.4 Interfaces
Related to the question of governance is the question of interfaces: how do different classes of stakeholders interact with the recommender systems? There is a great deal of study of consumer-side recommendation interfaces, and a wide variety of interface designs for end users to generate and interact with recommendations. Recommender systems interfaces for other stakeholders do exist, but are rarely the subject of published research. For example, YouTube provides a set of tools within their YouTube Studio application666studio.youtube.com to enable video creators to see some information about the viewership of their videos, but there are no detailed analytics about how the recommender system is handling their content or ways to interact with the recommender system itself.
The adversarial considerations noted above have no doubt deterred recommender system platforms from offering the kind of transparency into recommender system operations that other stakeholders might find useful. As a result, this is a highly underexplored aspect of multistakeholder recommender systems. Except for a few recent qualitative studies [8, 66], there is relatively little knowledge about provider-side experiences with recommender system interfaces.
5.5 Evaluation Design
The most widely used offline evaluation methodologies in the recommender systems are focused on user-oriented metrics like accuracy. When other stakeholders are considered, for example, in research that studies provider-side outcomes, researchers usually use the same methodology but evaluate the outcomes with provider-oriented metrics. One could imagine alternatives tailored to particular stakeholders: e.g., ensuring that items sampled in the test data set cover all providers, but there is little to no research on such stakeholder-specific evaluations.
Knijnenburg et al. [35] present a well-developed methodology for conducting user studies and interpreting them in terms of user experience. Such metrics might be exactly what is needed to understand different consumer-side aspects of a recommender system. There is no comparable methodology for understanding provider-side experiences of recommendation. It would only make sense to conduct user experience evaluation if an interface for providers exists, so this research area is downstream from the development of such interfaces.
5.6 Interactive / Conversational Recommendation
As of today, users are accustomed to one-shot static recommendations. Nevertheless, interactive/conversational systems are emerging as a technology that will likely change the nature of user interactions with recommender systems. The final outcome of a conversational session depends on the way the interaction is conducted by both parties: the user (consumer) and the system (that may behave on behalf of the provider). In a multistakeholder scenario, interaction is part of the overall recommendation process, and it is driven by the goals of the two actors involved in the conversation. In fact, depending on the conversation/interaction strategies, the final recommendation can be completely different and push towards the satisfaction of different goals of the involved stakeholders [32]. As a final observation, the interactive process itself may affect the satisfaction of some of the stakeholders’ goals. Among others, we may cite the number of interactions to get the final recommendation [11] or the seamless perception of the interactive process [44], but these are solely consumer-side metrics. There is little development of (for example) system-oriented metrics for conversational recommendation.
@@ Acknowledgments
The financial support by the Austrian Federal Ministry of Labour and Economy, the National Foundation for Research, Technology and Development and the Christian Doppler Research Association is gratefully acknowledged.
The work by Toine Bogers was supported by the FairMatch project (Innovation Fund Denmark grant number 3195-00003B). The work of Toine Bogers was also supported by the Pioneer Centre for AI (DNRF grant number P1).

### v2

@@ 4.1 Music Streaming
As noted above, measuring audience development can have a long time scale, so a short term proxy for this quality could be to measure to what extent an artist’s music is being recommended to receptive users. There are multiple ways to determine if a user is receptive [^6] For example, did the user listen to a second song by the artist, add their songs to a playlist, etc.?, but for the sake of example, let us assume that we can measure the number $n$ of non-audience listeners (that is, $u\notin A_{r}$) who were recommended a song by $r$ and then listened to the entire song. Given that different musicians have very different numbers of fans, it might make sense to normalize by the size of the artist’s existing audience $A_{r}$: $m_{r}=n/|A_{r}|$.
As a metric shared with individual providers, a low score on $m_{r}$ might raise concerns for the artist relative to the recommender system. It would mean that few new listeners are being introduced to their music. For a superstar, this might not be an issue: many people know their music already, but for an emerging artist, it could indicate that the recommender is not as helpful as it might be. A higher $m_{r}$ score does not necessarily mean that their audience is growing, but it does mean that the recommender system is introducing their music to potential new fans. From the system stakeholder point of view, this score could also be aggregated across all providers to understand audience building across the platform’s stable of artists. Its distribution might also be relevant in terms of fairness: are some types of artists better able to build audiences on the platform than others?
@@ 4.2 Education
Figure 4: Stakeholder configuration for the education example
@@ 4.3 Human Resources
Figure 5: Stakeholder configuration for the human resources example
@@
5 Supporting Stakeholder Insights and Control
A holistic approach to understanding and supporting the perspectives and needs of multiple stakeholders involves a range of challenges that go beyond what we have discussed so far. In the previous sections, we have described general properties of multistakeholder recommendation, and methodological approaches to developing relevant metrics, and outlined three hypothetical examples of metric development targeted to different classes of stakeholders. While several of these activities will require close cooperation and coordination with the different stakeholders, they do not yet address the question of how stakeholders can be empowered to gain insights relevant to them during system operation or even control some of the system’s functionalities. While these questions have increasingly been investigated in recent years for consumers as the main stakeholder class, very little research has yet addressed how other stakeholder classes might be supported in accessing and understanding information about the recommender system that serves their specific needs. Furthermore, the question arises whether, and to what extent stakeholders should be enabled to exert influence on the recommendation process itself. In this section, we discuss some salient challenges: Transparency & Explainability, Strategic / Adversarial Considerations, Interfaces and Interactive Control, and Governance.
5.1 Transparency and Explainability
@@
However, the types of evaluations that we discuss are different in that they may be of interest to parties who normally have no access to the workings of the recommender system. For example, the musical artists in our streaming example typically have very little insight into how the recommender system treats their content. While some major commercial music streaming platforms provide quite detailed analytics data for artists (see, e.g. Spotify  [^9] https://artists.spotify.com/analytics), the specific influence of the recommender system with respect to audience building or fairness is typically not exposed to the artists. However, we noted earlier that a musical artist might seek to understand to what extent the recommender system is helping them build an audience. Such metrics could be shared with artists as a form of explanation to help them understand, for example, how frequently their music is recommended in comparison to other artists (possibly with a similar genre and style) .
Explanations in a multistakeholder context also bring challenges different from explanations targeted only toward consumers. Firstly, different stakeholder groups have different explanatory needs that need to be identified. In the aforementioned example, the artist, their listeners, or the music label have quite different explanatory needs. This leads to the question of presenting different explanations to different stakeholders, or even personalizing them. For example, should we explain item recommendations to individual users and audience building to artists separately? When these concerns are integrated in an overall system objective, an additional type of explanation might be one that explains how we resolve the exposure of artists relative to how we weigh user preferences? In other words, if the requirement is to find and generate such a general explanation, then we need systems that can generate a meta-explanation on how tensions were resolved. A particular tension has already been observed where consumers (depending on the context of recommendation) may prefer less transparency in explanations if it gives better privacy [88, 121]. Members of groups preferred not to disclose sensitive information to other group members, and consumers of advertisements did not want certain information to be used in explanations (or as a basis of recommendations!). We do not attempt to answer the question of how to generate these kinds of explanations (as this is out of scope), but note that provider-side transparency, let alone generalized explanations, are very little studied in the context of multistakeholder recommendation.
@@ 5.2 Strategic / Adversarial Considerations
One likely reason that multistakeholder transparency has been little pursued is the concern that such a facility might be used to enable undesirable adversarial behavior. A web search for the term “YouTube algorithm” yields thousands of hits from search engine optimization (SEO) firms and others advising creators about how to bend the algorithm to their will. Additional information given to providers may enhance their ability to manipulate the algorithm in ways that are not necessarily beneficial to recommendation consumers or the platform. An open research question is how to offer provider-side disclosure in a way that limits adversarial opportunities. Where it is not possible to provide such disclosure, platforms may still seek to measure provider-side properties of their systems, but only for internal consumption.
5.3 Interfaces and Interactive Control
When translating transparency and explanation goals into concrete system designs, the question arises of how different classes of stakeholders can interact with the recommender system. There is a great deal of study of consumer-side recommendation interfaces, and a wide variety of interface designs for presenting recommendations, and for generating and interacting with recommendations. Recommender systems interfaces for other stakeholders do exist but are rarely the subject of published research. For example, YouTube provides a set of tools within their YouTube Studio application101010studio.youtube.com to enable video creators to see some information about the viewership of their videos, but there are no detailed analytics about how the recommender system is handling their content or ways to interact with the recommender system itself. Various suppliers of recommender system software or services also provide functionality for monitoring the recommendation process, mostly in the form of GUI-style dashboards. Depending on the platform, these dashboards present different types of analytic data, such as the number of personalized and non-personalized recommendations, the number of user interactions, or the temporal development of such data. However, the main target of monitoring dashboards are platform providers. Upstream or downstream stakeholders are generally not accommodated. Moreover, no specific evaluation approaches have become available so far for non-consumer stakeholders.
Today, consumers are used to static one-shot recommendations that cannot be influenced by them. In contrast, interactive and conversational systems are emerging as methods that can provide users with different levels of control over the recommendation process, turning the one-shot process into a multi-turn dialog ([56]. Examples of interactive methods include, for example, selecting or weighting the input data used by the algorithm, changing how strongly different item features influence the recommendation outcome, or interactively critiquing sample items. As we have noted, system and provider objectives may be quite different from those of consumers. We may expect that interactive recommendation techniques may strongly privilege the consumer side of the recommendation interaction; the impacts on other stakeholders of these technologies is still a matter for future research. Although methods for interacting with recommenders have mostly been GUI-based, the recent advances in language technologies have made text- or speech-based conversational approaches very promising. This holds both for controlling the recommendation process and for providing explanatory information [49]. However, conversational techniques have thus far also only been directed at consumers, but applying their capabilities to other stakeholder groups would provide these with flexible means to query or control the system according to their specific needs.
In parallel with lack of attention on interfaces and interaction for non-consumer stakeholders, specific evaluation metrics and methods are rare or missing, although the interactive process itself may strongly affect the satisfaction of some of the stakeholders’ goals. Existing interaction-related metrics such as the number of interactions to get the final recommendation [23] or the seamless perception of the interactive process [73], solely address the consumer side. There is also little development of user- or system-oriented metrics specifically for conversational recommendation techniques. As a result, interaction is an underexplored aspect of multistakeholder recommender systems. Except for a few recent qualitative studies [19, 102], there is relatively little knowledge about techniques and user experiences with recommender system interfaces for stakeholders along the entire value chain.
@@ 5.4 Governance
6 Conclusions
Multistakeholder recommender systems account for the impacts and preferences of multiple groups of individuals, not just the end users receiving recommendations. This article focuses on the evaluation of such multistakeholder recommender systems. We provide examples from three domains (music, HR, and education). Additionally, we discuss how to move from theoretical principles to practical implementation. We bring attention to the different aspects involved—from the range of stakeholders involved (including but not limited to producers and consumers) to the values and specific goals of each relevant stakeholder. Next, we discuss evaluation methodologies and metric selection. Finally, we outline several open areas for future research, as seen from the lens of multistakeholder evaluation.
The structure of the paper follows the steps necessary to evaluate a methodology for software application development through: defining relevant features; scoring the features; and providing guidelines for applying the evaluation framework [58]. Overall, with this work, we aimed to outline the components that cannot be overlooked when undertaking a multistakeholder evaluation of recommender systems. As we have emphasized throughout the narrative, it cannot be denied that each use case is unique, influenced by the characteristics of the primary recommender system, the environment in which it is deployed, and the (direct and indirect) users it is designed to serve. Nevertheless, the components we discuss, along with the focus on the practical implementation of key areas—from identifying stakeholders and eliciting their views on recommender systems to defining and measuring how well the envisioned values for a recommender are upheld—should serve as guidance for researchers and practitioners. This understanding is essential for navigating the complex task of evaluating recommender systems that satisfy multiple stakeholders while considering the necessary trade-offs and compromises that need to be made along the way.
@@ Acknowledgments
The financial support by the Austrian Federal Ministry of Labour and Economy, the National Foundation for Research, Technology and Development and the Christian Doppler Research Association for research by Julia Neidhardt is gratefully acknowledged.
The work of Robin Burke was supported by the National Science Foundation under award IIS-2107577.
The work by Toine Bogers was supported by the FairMatch project (Innovation Fund Denmark grant number 3195-00003B) and by the Pioneer Centre for AI (DNRF grant number P1).